# Stability as a Liability: Systematic Breakdown of Linguistic Structure in LLMs

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
However, stability is not a monolithic concept. Existing methods primarily stabilize task-aligned objectives, such as reward signals(Learning, ), preference models(Gumen and Savochkin, 2013), or supervised losses(Akhtar et al., 2024). In contrast, little attention has been paid to the stability of internal generation dynamics, including token-level distributions, entropy evolution, and termination behavior. Whether stabilizing these internal dynamics is always beneficial remains underexplored.
@@
We conduct a comprehensive empirical analysis using entropy trajectories, effective rank measurements, length distributions, and cross-seed consistency tests. Our results show that lower training loss and increased stability do not necessarily imply improved or even preserved generative capacity.
@@
    We identify a previously underexplored failure mode in neural language model training, where optimization remains stable under standard metrics while semantic generation degrades systematically, revealing a disconnect between training stability and generative quality.
@@
    We identify a previously underexplored failure mode in neural language model training, where optimization remains stable under standard metrics while semantic generation degrades systematically, revealing a disconnect between training stability and generative quality.
@@
    We provide a theoretical interpretation of this phenomenon by analyzing how excessive stabilization constrains the effective degrees of freedom during training, leading to reduced expressive capacity at generation time despite apparent convergence.
@@ 3 Problem Setting
While Definition 1 characterizes stability purely in terms of optimization dynamics in parameter space, our interest lies in how such stability constraints propagate to the induced generation distribution $P_{\theta}$. In autoregressive language models, parameters and distributions are coupled through the mapping $\theta\mapsto P_{\theta}$, and restrictions on the parameter trajectory implicitly restrict the set of attainable output distributions.
@@ 4.3 Impact of Approximate KL Optimality
By Lemma 1, let $\theta^{*}$ be a stationary point of the regularized MLE objective
@@
Taking the Euclidean norm on both sides and using the asymptotic parameter boundedness from Lemma 1, $\|\theta^{*}\|\leq M<\infty$, we have
@@ 5 Mode Collapse Induced by Training Stability
Lemma 3 shows that any stationary point $\theta$ approximately minimizes the forward KL divergence up to $O(\lambda)$. To analyze the implications of this property for mode collapse, we formalize the assumptions on the training process and the data distribution as follows.
@@
The training process is stable in the sense of Definition 1: the parameter sequence $\{\theta_{t}\}$ is asymptotically bounded, the gradients vanish asymptotically, and the loss $\mathcal{L}(\theta)$ is $\beta$-smooth.
@@
By Lemma 3, $\theta^{*}$ is an approximate minimizer of the forward KL divergence:
@@ 6.1 Experimental Setup
Figure 1: Framework for studying mode collapse: feedback network with BARRA module.
To empirically validate Theorem 5.5, we construct a closed-loop control network(Zheng et al., 2024) tailored to probe stability-induced mode collapse. The network implements a first-order negative feedback system with discrete-time dynamics
@@
where $\alpha>0$ is the step size, $f$ encodes the feedback function, and $\eta_{t}$ is stochastic noise. Learning is stabilized via our PSU algorithm, which biases parameter trajectories toward asymptotic boundedness and vanishing gradients, satisfying the stable training assumptions of Definition 1. This configuration ensures the system dynamics are tractable while allowing direct observation of emergent behavior under stable training.
To systematically explore the network’s output variability, we integrate two stochastic modules: The Dynamic MLP Unit (DMU) randomly drops activations at each step to regulate learning intensity and enforce sparsity of active modes, while the Dynamic Transformer Unit (DTU) injects stochastic perturbations into parameter updates, simulating variability in learning paths. Together, these modules maintain controlled exploration within the stable training regime, enabling the empirical distribution of outputs to concentrate on specific modes. The resulting probability concentration is monitored over time to quantify mode collapse. Figure 4 provides a schematic overview of the system. Our detailed impletation and analysis of experiment can be seen in Appendix. A and B. Codes are available at https://anonymous.4open.science/r/Stability-as-a-Liability-Systematic-Breakdown-of-Linguistic-Structure-in-LLMs-99F7.
@@ 6.3 Ablation Studies
We conduct ablation studies on BARRA’s DMU and DTU to validate that stabilization induces mode collapse via forward KL minimization and entropy reduction, as shown in Fig.2. Removing each component and comparing to the full BARRA baseline we quantify how the two components synergistically amplify stabilization effects worsening the entropy-loss trajectory driving mode collapse. All configurations share identical setup ensuring performance differences reflect BARRA’s stabilization impact on the tradeoff between optimization stability and generative degradation.
Figure 2: Ablation Studies
@@
These results validate our theory that stability is a liability. Differential performance at moderate $\alpha$ links BARRA’s components to forward KL minimization and entropy reduction, tying stabilization to generative degradation. Conditional activation of BARRA components underscores that mode collapse is stability-dependent and rooted in MLE properties. Accelerated collapse in the baseline at extreme $\alpha$ confirms stabilization optimizes optimization dynamics at the cost of generative expressivity. Collectively, these findings establish that the BARRA framework validates our core thesis: stability induces systematic generative degradation, highlighting the tension between optimization reliability and generative quality.
@@ 6.4 Loss-Entropy Trajectory and Mode Collapse Vulnerability
To further validate how optimization dynamics drive mode collapse we analyze the joint trajectory of training loss and generative entropy, shown in Fig.3. This relationship reveals the inherent vulnerability of low-loss regimes to mode collapse.
Figure 3: Relationship between Loss and Entropy in LLM Training
@@ 7 Conclusion
In this work we establish a rigorous connection between stable training trajectories and mode collapse in generative models. We formalize conditions where stationary points of stabilized MLE objectives concentrate probability mass on a small subset of empirical support and verify these predictions in a controlled closed loop feedback system with stochastic exploration modules. Our analysis clarifies the fundamental tension between stability and diversity in high-dimensional learning providing a principled framework to anticipate measure and potentially mitigate mode collapse. Theoretical guarantees and empirical validation highlight the importance of parameter dynamics in designing robust generative systems. This work lays the foundation for future explorations into balancing stability expressivity and controlled stochasticity in generative modeling.
@@ Acknowledgements
Appendix A Definitions and Notations
@@ B.3.1 Primary Algorithms for MLP
For an $L$-layer MLP, node-level and edge-level dropout are asymptotically equivalent in expectation and normalized variance for parameter updates and final output as $\min_{\ell=1}^{L}n_{\ell}\to\infty$.
@@ B.4 MLP-Transformer Consistency
We choose dimensions in X to calculate and change $a_{2}$ to be $a_{3}$. $a_{3}$ is smaller than $a_{2}$. According to Theorem 3.1 and Theorem 3.3, we can choose each element to save for proposition $a_{3}/a_{2}$.

### v2

@@ Abstract
Large Language Models, Training Stability, Generative Entropy, Linguistic Structure
@@ 1 Introduction
However, stability is not a monolithic concept. Existing methods primarily stabilize task-aligned objectives, such as reward signals (Zhang et al., 2025), preference models (Gumen and Savochkin, 2013), or supervised losses (Akhtar et al., 2024). In contrast, little attention has been paid to the stability of internal generation dynamics, including token-level distributions, entropy evolution, and termination behavior. Whether stabilizing these internal dynamics is always beneficial remains underexplored.
@@
We conduct a comprehensive empirical analysis using entropy trajectories, effective rank measurements, length distributions, and cross-seed consistency tests. Our results show that lower training loss and increased stability do not necessarily imply improved or even preserved generative capacity. Codes are available at https://anonymous.4open.science/r/Stability-as-a-Liability-Systematic-Breakdown-of-Linguistic-Structure-in-LLMs-99F7.
@@
    Identify a failure mode in large language model training: stable optimization under standard metrics coexists with systematic degradation of semantic generation performance, confirming an intrinsic disconnect between training stability and generative quality.
@@
    Verify the robustness of this failure mode across architectures, parameter scales and random seeds via a controlled closed-loop training framework, ruling out interference factors like training failure, hyperparameter selection or insufficient model capacity.
@@
    Propose a theoretical interpretation for excessive stabilization in MLE training, showing it restricts effective degrees of freedom and reduces generative expressivity even with parameter convergence, quantifying stability-mode collapse correlation.
@@ 3 Problem Setting
While Definition 3.1 characterizes stability purely in terms of optimization dynamics in parameter space, our interest lies in how such stability constraints propagate to the induced generation distribution $P_{\theta}$. In autoregressive language models, parameters and distributions are coupled through the mapping $\theta\mapsto P_{\theta}$, and restrictions on the parameter trajectory implicitly restrict the set of attainable output distributions.
@@ 4.3 Impact of Approximate KL Optimality
By Lemma 4.1, let $\theta^{*}$ be a stationary point of the regularized MLE objective
@@
Taking the Euclidean norm on both sides and using the asymptotic parameter boundedness from Lemma 4.1, $\|\theta^{*}\|\leqslant M<\infty$, we have
@@ 5 Mode Collapse Induced by Training Stability
Lemma 4.3 shows that any stationary point $\theta$ approximately minimizes the forward KL divergence up to $O(\lambda)$. To analyze the implications of this property for mode collapse, we formalize the assumptions on the training process and the data distribution as follows.
@@
The training process is stable in the sense of the following condition: the parameter sequence $\{\theta_{t}\}$ is asymptotically bounded, the gradients vanish asymptotically, and the loss $\mathcal{L}(\theta)$ is $\beta$-smooth.
@@
By Lemma 4.3, $\theta^{*}$ is an approximate minimizer of the forward KL divergence:
@@ 6.1 Experimental Setup
Figure 1: Framework for studying mode collapse: feedback network with BARRA module. This closed-loop control system stabilizes large language model generation dynamics, integrating stochastic modules to regulate learning intensity and inject controlled variability. It isolates stability-induced mode collapse by decoupling dynamics stabilization from semantic optimization, capturing key degenerative behaviors like low-entropy outputs and repetitive token sequences.
To empirically validate Theorem 5.5, we construct a closed-loop control network tailored to probe stability-induced mode collapse. The network implements a first-order negative feedback system with discrete-time dynamics
@@
where $\alpha>0$ is the step size, $f$ encodes the feedback function, and $\eta_{t}$ is stochastic noise. Learning is stabilized via our PSU algorithm, which biases parameter trajectories toward asymptotic boundedness and vanishing gradients, satisfying the stable training assumptions of Assumption 5.1. This configuration ensures the system dynamics are tractable while allowing direct observation of emergent behavior under stable training.
To systematically explore the network’s output variability, we integrate two stochastic modules: The Dynamic MLP Unit (DMU) randomly drops activations at each step to regulate learning intensity and enforce sparsity of active modes, while the Dynamic Transformer Unit (DTU) injects stochastic perturbations into parameter updates, simulating variability in learning paths. Together, these modules maintain controlled exploration within the stable training regime, enabling the empirical distribution of outputs to concentrate on specific modes. The resulting probability concentration is monitored over time to quantify mode collapse. Figure 1 provides a schematic overview of the system. Our detailed impletation and analysis of experiment can be seen in Appendix A and  B.
@@ 6.3 Ablation Studies
Figure 2: Ablation Studies of BARRA’s DMU/DTU Components. High-frequency word proportion across $\alpha$ values quantifies synergistic stabilization effects on mode collapse via KL minimization and entropy reduction. Results show DMU and DTU jointly intensify mode concentration underperforming single-module ablations.
We conduct ablation studies on BARRA’s DMU and DTU to validate that stabilization induces mode collapse via forward KL minimization and entropy reduction, as shown in Figure. 2. Removing each component and comparing to the full BARRA baseline we quantify how the two components synergistically amplify stabilization effects worsening the entropy-loss trajectory driving mode collapse. All configurations share identical setup ensuring performance differences reflect BARRA’s stabilization impact on the tradeoff between optimization stability and generative degradation.
@@
These results validate our theory that stability is a liability. Differential performance at moderate $\alpha$ links BARRA’s components to forward KL minimization and entropy reduction, tying stabilization to generative degradation. Conditional activation of BARRA components underscores that mode collapse is stability-dependent and rooted in MLE properties. Accelerated collapse in the baseline at extreme $\alpha$ confirms stabilization optimizes optimization dynamics at the cost of generative expressivity.
@@ 6.4 Loss-Entropy Trajectory and Mode Collapse Vulnerability
Figure 3: Joint Trajectory of Training Loss and Generative Entropy in LLM Training. Color intensity marks point density, revealing the tradeoff between loss minimization and generative diversity across training steps. High-loss stages show sharp entropy drops while low-loss regimes yield marginal entropy reductions, reflecting pathological probability concentration.
To further validate how optimization dynamics drive mode collapse we analyze the joint trajectory of training loss and generative entropy, shown in Figure.3. This relationship reveals the inherent vulnerability of low-loss regimes to mode collapse.
@@ 7 Conclusion
In this work we establish a rigorous connection between stable training trajectories and mode collapse in generative models. We formalize conditions where stationary points of stabilized MLE objectives concentrate probability mass on a small subset of empirical support and verify these predictions in a controlled closed-loop feedback system with stochastic exploration modules. Our analysis clarifies the fundamental tension between stability and diversity in high-dimensional learning providing a principled framework to anticipate measure and potentially mitigate mode collapse. Theoretical guarantees and empirical validation across architectures and parameter scales highlight the importance of parameter dynamics in designing robust generative systems. This work lays the foundation for future explorations into balancing stability expressivity and controlled stochasticity in generative modeling.
@@ Acknowledgements
Appendix A Definition and Notations
@@ B.3.1 Primary Algorithms for MLP
For an $L$-layer MLP satisfying Assumptions 1–3, node-level (DProb/DFix) and edge-level (DMU2-Prob/DMU2-Fix) dropout are asymptotically equivalent in expectation and normalized variance for parameter updates and final output as $\min_{\ell=1}^{L}n_{\ell}\to\infty$.
@@ B.4 MLP-Transformer Consistency
We choose dimensions in X to calculate and change $a_{2}$ to be $a_{3}$. $a_{3}$ is smaller than $a_{2}$. We can choose each element to save for proposition $a_{3}/a_{2}$.